# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v2

@@ 7 EXPERIMENTS
We conduct a series of experiments, each having a different source of auxiliary data, to test the proposed DR ARO models using empirical and auxiliary datasets. We use the following abbreviations, where ‘solution’ refers to the optimal $\bm{\beta}$ to make decisions:
@@
    DRO+ARO+Aux: Solution of 16 (i.e., relaxation of Inter-ARO that revises DR-ARO and intersects its ambiguity set with a Wasserstein ball built using auxiliary data);
Recall that DRO+ARO and DRO+ARO+Aux are the DR models that we propose. Note also that, ERM, ARO, and DRO+ARO are oblivious to auxiliary data. All Wasserstein radii of DR models, and the weight parameters of ARO+Aux are cross-validated from the same grids. All experiments are conducted in Julia (Bezanson et al.,, 2014) and executed on Intel Xeon 2.66GHz processors with 16GB memory in single-core mode. We use MOSEK’s exponential cone optimizer to solve all problems. Implementation details are in the appendices.
@@ 7.1 UCI Datasets (Auxiliary Data is Synthetically Generated)
We compare the out-of-sample error rates of each method on 10 UCI datasets for binary classification (Kelly et al.,, 2023). For each dataset, we run 10 simulations as follows: (i) Select $40\%$ of the data as a test set ($N_{\mathrm{te}}\propto 0.4$); (ii) Sample $25\%$ of the remaining to form a training set ($N\propto 0.6\cdot 0.25$); (iii) The rest ($\widehat{N}\propto 0.6\cdot 0.75$) is used to fit a synthetic generator Gaussian Copula from the SDV package (Patki et al.,, 2016), which is then used to generate auxiliary data. The mean errors on the test set are reported in Table 1 for $\ell_{2}$-attacks of strength $\alpha=0.05$. The best error is always achieved by DRO+ARO+Aux, followed by DRO+ARO, DRO+Aux, ARO, ERM, respectively. In our appendices, we report similar results for attack strengths $\alpha\in\{0,0.05,0.2\}$, and share data preprocessing details and standard deviations of out-of-sample errors.
@@ 7.2 MNIST/EMNIST Datasets (Auxiliary Data is Out-of-Domain)
We use the MNIST digits dataset (LeCun et al.,, 1998) to classify whether a digit is 1 or 7. For an auxiliary dataset, we use the larger EMNIST digits dataset (Cohen et al.,, 2017), whose authors summarize that this dataset has additional samples collected from a different group of individuals (high school students). Since EMNIST digits include MNIST digits, we remove the latter from the EMNIST dataset. We simulate the following 25 times: (i) Sample 1,000 instances from the MNIST dataset as a training set; (ii) The remaining instances in the MNIST dataset are our test set; (iii) Sample 1,000 instances from the EMNIST dataset as an auxiliary dataset. Table 2 reports the mean out-of-sample errors in various adversarial attack regimes. The results are analogous to the UCI experiments.
@@ 7.3 Artificial Experiments (Auxiliary Data is Perturbed)
We generate empirical and auxiliary datasets by controlling their data-generating distributions (more details in the appendices). We simulate $25$ cases, each with $N=100$ training, $\widehat{N}=200$ auxiliary, and $N_{\mathrm{te}}=10,000$ test instances and $n=100$ features. The performance of benchmark models with varying $\ell_{2}$-attacks is available in Figure 2 (left). ERM provides the worst performance, followed by ARO. The relationship between DRO+ARO and ARO+Aux is not monotonic: the latter works better in larger attack regimes, conforming to the robust overfitting phenomenon. Finally, Adv+DRO+Aux always performs the best. We conduct a similar simulation for datasets with $n=100$, and gradually increase $N=\widehat{N}$ to report median ($50\%\pm 15\%$ quantiles shaded) runtimes of each method (cf. Figure 2, right). The fastest methods is ARO, followed by ERM, ARO+Aux, DRO+ARO, and DRO+ARO+Aux. The slowest is DRO+ARO+Aux, but the runtime scales graciously.
@@ 8 CONCLUSIONS
We formulate the distributionally robust counterpart of adversarially robust logistic regression. Additionally, we demonstrate how to effectively utilize appropriately curated auxiliary data by intersecting Wasserstein balls. We illustrate the superiority of the proposed approach in terms of out-of-sample performance and confirm its scalability in practical settings.
@@
It would be natural to extend our results to more loss functions as is typical for theoretical DRO studies stemming from logistic regression. Moreover, recent breakthroughs in the area of foundation models naturally pose the question of whether our ideas in this work apply to these models. For example, Ye et al., (2022) use pre-trained language model (PLM) to generate synthetic pairs of text sequences and labels which are then used to train downstream models. It would be interesting to adapt our ideas to the text domain to explore robustness in the presence of two PLMs.
Supplementary Materials
9 NOTATION
@@
Throughout the paper, bold lower case letters denote vectors, while standard lower case letters are reserved for scalars. A generic data instance is modeled as $\bm{\xi}=(\bm{x},y)\in\Xi:=\mathbb{R}^{n}\times\{-1,+1\}$. For any $p>0$, $\lVert\bm{x}\rVert_{p}$ denotes the rational norm $\left(\sum_{i=1}^{n}\lvert x_{i}\rvert^{p}\right)^{1/p}$ and $\lVert\bm{x}\rVert_{p^{\star}}$ is its dual norm where $11/p+1/p^{\star}=$ 1 with the convention of $/1+1/\infty=1$ 1. The set of probability distributions supported on $\Xi$ is denoted by $\mathcal{P}(\Xi)$. The Dirac measure supported on $\bm{\xi}$ is denoted by $\delta_{\bm{\xi}}$. The logloss is defined as $\ell_{\bm{\beta}}(\bm{x},y)=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}))$ and its associated univariate loss is $L(z)=\log(1+\exp(-z))$ so that $L(y\cdot\bm{\beta}^{\top}\bm{x})=\ell_{\bm{\beta}}(\bm{x},y)$. The exponential cone is denoted by $\mathcal{K}_{\exp}=\mathrm{cl}(\{\bm{\omega}\in\mathbb{R}^{3}:\omega_{1}\geq\omega_{2}\cdot\exp(\omega_{3}/\omega_{2}),\ \omega_{1}>0,\ \omega_{2}>0\})$ where $\mathrm{cl}$ is the closure operator. The Lipschitz modulus of a univariate function $f$ is defined as $\mathrm{Lip}(f):=\sup_{z,z^{\prime}\in\mathbb{R}}\left\{\lvert f(z)-f(z^{\prime})\rvert/\lvert z-z^{\prime}\rvert:z\neq z\right\}$ whereas its effective domain is $\mathrm{dom}(f)=\{z:f(z)<+\infty\}$. For a function $f:\mathbb{R}^{n}\mapsto\mathbb{R}$, its convex conjugate is $f^{*}(\bm{z})=\sup_{\bm{x}\in\mathbb{R}^{n}}\bm{z}^{\top}\bm{x}-f(\bm{x})$. We reserve $\alpha\geq 0$ for the radii of the norms of adversarial attacks on the features and $\varepsilon\geq 0$ for the radii of distributional ambiguity sets.
10 MISSING PROOFS
10.1 Proof of Observation 1
@@
For any $\bm{\beta}\in\mathbb{R}^{n}$, with standard robust optimization arguments (Ben-Tal et al.,, 2009; Bertsimas and Den Hertog,, 2022), we can show that
@@
where the first step follows from the definition of logloss, the second step follows from the fact that $\log$ and $\exp$ are increasing functions, the third step takes the constant terms out of the $\sup$ problem and exploits the fact that the optimal solution of maximizing a linear function will be at an extreme point of the $\ell_{p}$ ball, the fourth step uses the definition of dual norm, and finally the redundant $-y\in\{-1,+1\}$ is omitted from the dual norm. We can therefore define the adversarial loss $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y):=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ where $\alpha$ models the strength of the adversary, $\bm{\beta}$ is the decision vector, and $(\bm{x},y)$ is an instance. Replacing ${\sup_{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}$ in DR-ARO with $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$ concludes the equivalence of the optimization problem.
@@ 10.2 Proof of Corollary 1
Since the univariate loss $L^{\alpha}(z):=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ satisfying the identity $L^{\alpha}(\langle y\cdot\bm{x},\bm{\beta}\rangle)=\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$ is Lipschitz continuous, Theorem 14 (ii) of Shafieezadeh-Abadeh et al., (2019) is immediately applicable. We can therefore rewrite (24) as:
@@
10.3 Proof of Proposition 1
We prove Proposition 1 by constructing the optimization problem in its statement. We will thus dualize the inner $\sup$ problem of Inter-ARO for fixed $\bm{\beta}$. To this end, we present a sequence of reformulations to the inner problem and then exploit strong semi-infinite duality.
@@ 10.3 Proof of Proposition 1
The following semi-infinite optimization problem, obtained by standard algebraic duality, is a strong dual to the above problem since $\varepsilon,\widehat{\varepsilon}>0$ (Shapiro,, 2001).
@@
$\begin{array}[]{cll}\displaystyle\underset{\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}}}{\mathrm{minimize}}&\displaystyle\dfrac{1}{N}\left[N\varepsilon\lambda+\widehat{N}\widehat{\varepsilon}\widehat{\lambda}+\sum_{i=1}^{N}s_{i}+\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\right]&\\[14.22636pt] \mathrm{subject\;to}&\displaystyle\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\{\ell^{\alpha}_{\bm{\beta}}(\bm{x},+1)-\lambda\lVert\bm{x}^{i}-\bm{x}\rVert_{q}-\dfrac{\widehat{N}}{N}\widehat{\lambda}\lVert\widehat{\bm{x}}^{j}-\bm{x}\rVert_{q}\}\leq\\[8.53581pt] &{\quad s_{i}+\kappa\dfrac{1-y^{i}}{2}\lambda+\dfrac{\widehat{N}}{N}\cdot\left[\widehat{s}_{j}+\kappa\dfrac{1-\widehat{y}^{j}}{2}\widehat{\lambda}\right]}&\forall i\in[N],\ \forall j\in[\widehat{N}]\\[14.22636pt] &\displaystyle\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\{\ell^{\alpha}_{\bm{\beta}}(\bm{x},-1)-\lambda\lVert\bm{x}^{i}-\bm{x}\rVert_{q}-\dfrac{\widehat{N}}{N}\widehat{\lambda}\lVert\widehat{\bm{x}}^{j}-\bm{x}\rVert_{q}\}\leq\\[8.53581pt] &{\quad s_{i}+\kappa\dfrac{1+y^{i}}{2}\lambda+\dfrac{\widehat{N}}{N}\cdot\left[\widehat{s}_{j}+\kappa\dfrac{1+\widehat{y}^{j}}{2}\widehat{\lambda}\right]}&\forall i\in[N],\ \forall j\in[\widehat{N}]\\[14.22636pt] &\lambda\geq 0,\ \widehat{\lambda}\geq 0,\ \bm{s}\in\mathbb{R}^{N}_{+},\ \widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+}\end{array}$
@@
$\begin{array}[]{lll}&\displaystyle\int_{\bm{x}\in\mathbb{R}^{n}}\mathbb{Q}_{+1}^{i}(\mathop{}\!\mathrm{d}\bm{x})+\mathbb{Q}_{-1}^{i}(\mathop{}\!\mathrm{d}\bm{x})\leq 1&\forall i\in[N]\\[14.22636pt] &\displaystyle\int_{\bm{x}\in\mathbb{R}^{n}}\widehat{\mathbb{Q}}_{+1}^{j}(\mathop{}\!\mathrm{d}\bm{x})+\widehat{\mathbb{Q}}_{-1}^{j}(\mathop{}\!\mathrm{d}\bm{x})\leq 1&\forall j\in[\widehat{N}]\end{array}$
@@
Combining all the $\sup$ constraints with the help of an an auxiliary parameter $l\in\{-1,1\}$ and replacing this problem with the inner problem of Inter-ARO concludes the proof.∎
10.4 Proof of Proposition 2
We first present a technical lemma that will allow us to rewrite a specific type of difference of convex functions (DC) maximization problem that appears in the constraints of Inter-ARO. Rewriting such DC maximization problems is one of the key steps in reformulating Wasserstein DRO problems, and our lemma is inspired from Shafieezadeh-Abadeh et al., (2019, Lemma 47)Shafieezadeh-Abadeh et al., (2023, Theorem 3.8), and Belbasi et al., (2023, Lemma 1) who reformulate maximizing the difference of a convex function and a norm. Our DRO problem Inter-ARO, however, comprises two ambiguity sets, hence the DC term that we investigate will be the difference between a convex function and the sum of two norms. This requires a new analysis and we will see that Inter-ARO is NP-hard due to this additional difficulty.

### v4

@@ 7 EXPERIMENTS
We conduct a series of experiments, each having a different source of auxiliary data, to test the proposed DR ARO models. We use the following abbreviations, where ‘solution’ refers to the optimal $\bm{\beta}$ to make decisions:
@@
    DRO+ARO+Aux: Solution of 21 (i.e., relaxation of Inter-ARO that intersects the ambiguity set of DR-ARO with an auxiliary Wasserstein ball);
All parameters are $5$-fold cross-validated from various grids. The Wasserstein radii use the grid $\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},0,1,2,5,10\}$, which is sufficient to ensure that the rule-of-thumb $\varepsilon=\mathcal{O}(1/\sqrt{N})$ is included around the center of this grid for all experiments conducted. To ensure that the intersections of Wasserstein balls are nonempty, we compute $\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$ once, and discard all combinations $(\varepsilon,\widehat{\varepsilon})$ with $\varepsilon+\widehat{\varepsilon}<\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$. The weight parameter $\omega$ of ARO+Aux is cross-validated from the grid $\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},0,1\}$. We fix the norms defining the feature-label metric and the adversarial attacks to $\ell_{1}$- and $\ell_{2}$-norms, respectively. The parameter $\kappa$ (cf. Definition 1) is cross-validated from the grid $\{1,\sqrt{n},n\}$, and since $n$ is the number of features, this grid includes cases where label uncertainty is equivalent to uncertainty of a single feature ($\kappa=1$), label uncertainty is equivalent to uncertainty of all features combined ($\kappa=n$), and an intermediary case ($\kappa=\sqrt{n}$). The case of ignoring label uncertainty ($\kappa=\infty$) is purposely not included in the grid, since one of the key reasons behind robust overfitting is that while ARO is equivalent to a distributionally robust model, the underlying ambiguity is only around the features (cf. our discussion in §2). All simulated adversarial attacks are worst-case $\ell_{p}$-attacks that are instance-wise at test time, and the experiments assume we know the strength $\alpha$ and norm $\ell_{p}$ of the adversarial attacks.
All experiments are conducted in Julia and executed on Intel Xeon 2.66GHz processors with 16GB memory in single-core mode. We use MOSEK’s exponential cone optimizer to solve all problems. To interpret the results accurately, recall that DRO+ARO and DRO+ARO+Aux are the DR models that we propose. Note also that ERM, ARO, and DRO+ARO do not utilize auxiliary data, while DRO+ARO+Aux and ARO+Aux have access to the same auxiliary datasets across all experiments (i.e., we do not sample different auxiliary distributions for different methods to ensure that our comparisons are made ceteris paribus). Moreover, while ARO does not have access to auxiliary data, one can interpret ARO+Aux as a generalization of ARO that also has access to auxiliary datasets since it takes a mixture (with mixture weight $\omega$) of the empirical dataset with the auxiliary dataset. For example, $\omega=1$ would simply revise ARO by appending the empirical dataset with the auxiliary dataset.
@@ 7.1 UCI Datasets (Auxiliary Data is Synthetically Generated)
We compare the out-of-sample error rates of each method on 10 UCI datasets for binary classification (Kelly et al., 2023). For each dataset, we run 10 simulations as follows: (i) Select $40\%$ of the data as a test set ($N_{\mathrm{te}}\propto 0.4$); (ii) Sample $25\%$ of the remaining to form a training set ($N\propto 0.6\cdot 0.25$); (iii) The rest ($\widehat{N}\propto 0.6\cdot 0.75$) is used to fit a synthetic generator Gaussian Copula from the SDV package (Patki et al., 2016), which is then used to generate auxiliary data. The mean errors on the test set are reported in Table 1 for $\ell_{2}$-attacks of strength $\alpha=0.05$. The best error is always achieved by DRO+ARO+Aux, followed by DRO+ARO, DRO+Aux, ARO, ERM, respectively. In our appendices, we report similar results for attack strengths $\alpha\in\{0,0.05,0.2\}$, and share data preprocessing details and standard deviations of out-of-sample errors.
@@ 7.2 MNIST/EMNIST Datasets (Auxiliary Data is Out-of-Domain)
We use the MNIST digits dataset (LeCun et al., 1998) to classify whether a digit is 1 or 7. For an auxiliary dataset, we use the larger EMNIST digits dataset (Cohen et al., 2017), whose authors summarize that this dataset has additional samples collected from a different group of individuals (high school students). Since EMNIST digits include MNIST digits, we remove the latter from the EMNIST dataset. We simulate the following 25 times: (i) Sample 1,000 instances from the MNIST dataset as a training set; (ii) The remaining instances in the MNIST dataset are our test set; (iii) Sample 1,000 instances from the EMNIST dataset as an auxiliary dataset. Table 2 reports the mean out-of-sample errors in various adversarial attack regimes. The results are analogous to the UCI experiments. Additionally, note that in the absence of adversarial attacks ($\alpha=0$), DRO+ARO coincides with the Wasserstein LR model of Shafieezadeh-Abadeh et al. (2015), and the results thus imply that even without adversarial attacks, we can improve the state-of-the-art DR model by revising its ambiguity set in light of auxiliary data.
@@ 7.3 Artificial Experiments (Auxiliary Data is Perturbed)
We generate empirical and auxiliary datasets by controlling their data-generating distributions (more details in the appendices). We simulate $25$ cases, each with $N=100$ training, $\widehat{N}=200$ auxiliary, and $N_{\mathrm{te}}=10,000$ test instances and $n=100$ features. The performance of benchmark models with varying attacks is available in Figure 2 (left). ERM provides the worst performance, followed by ARO. The relationship between DRO+ARO and ARO+Aux is not monotonic: the former works better in larger attack regimes, conforming to the robust overfitting phenomenon. Finally, Adv+DRO+Aux always performs the best. We conduct a similar simulation for datasets with $n=100$, and gradually increase $N=\widehat{N}$ to report median ($50\%\pm 15\%$ quantiles shaded) runtimes of each method (cf. Figure 2, right). The fastest methods is ARO, followed by ERM, ARO+Aux, DRO+ARO, and DRO+ARO+Aux. The slowest is DRO+ARO+Aux, but the runtime scales graciously.
@@ 8 CONCLUSIONS
We formulate the distributionally robust counterpart of adversarially robust LR. Additionally, we demonstrate how to effectively utilize appropriately curated auxiliary data by intersecting Wasserstein balls. We illustrate the superiority of the proposed approach in terms of out-of-sample performance and confirm its scalability in practical settings.
From a theoretical point of view, it would be natural to extend our work to more loss functions, as is typical for DRO studies stemming from LR. To be able to optimize 21 for very large-dimensional datasets, an interesting future work is to investigate first-order optimization methods that do not rely on off-the-shelf solvers. We also believe a cutting-plane method tailored for 21 can also help us scale this problem for large-dimensional problems, since we would avoid monolithically optimizing a problem with $\mathcal{O}(N\cdot\widehat{N})$ exponential cone constraints.
From a practical perspective, the ability to optimize 21 in high-dimensional settings would also enable fine-tuning the final layer of a pre-trained neural network for binary classification, since this corresponds to logistic regression under a sigmoid activation. In our image recognition experiments, we used the MNIST dataset, as EMNIST served as a natural choice for auxiliary data. Identifying a suitable auxiliary dataset for CIFAR (Krizhevsky et al., 2009) could similarly support new experimental directions.
@@
Finally, recent breakthroughs in foundation models naturally pose the question of whether our ideas in this work apply to these models. For example, Ye et al. (2022) use a pre-trained language model (PLM) to generate synthetic pairs of text sequences and labels which are then used to train downstream models. It would be interesting to adapt our ideas to the text domain to explore robustness in the presence of two PLMs.
Acknowledgements.
Aras Selvi’s work was done during an internship at JP Morgan AI Research. The authors gratefully acknowledge the detailed and constructive feedback provided by the anonymous reviewers and the anonymous area chair. Revising the paper has substantially improved the quality of this work.
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (“JP Morgan”) and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix A NOTATION
@@
Throughout the paper, bold lowercase letters denote vectors, while standard lowercase letters are reserved for scalars. A generic data instance is modeled as $\bm{\xi}=(\bm{x},y)\in\Xi:=\mathbb{R}^{n}\times\{-1,+1\}$. For any $p>0$, $\lVert\bm{x}\rVert_{p}$ denotes the $p$-norm $\left(\sum_{i=1}^{n}\lvert x_{i}\rvert^{p}\right)^{1/p}$ and $\lVert\bm{x}\rVert_{p^{\star}}$ is its dual norm where $11/p+1/p^{\star}=$ 1 with the convention of $/1+1/\infty=1$ 1. The set of probability distributions supported on $\Xi$ is denoted by $\mathcal{P}(\Xi)$. The Dirac measure supported on $\bm{\xi}$ is denoted by $\delta_{\bm{\xi}}$. The logloss is defined as $\ell_{\bm{\beta}}(\bm{x},y)=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}))$ and its associated univariate loss is $L(z)=\log(1+\exp(-z))$ so that $L(y\cdot\bm{\beta}^{\top}\bm{x})=\ell_{\bm{\beta}}(\bm{x},y)$. The exponential cone is denoted by $\mathcal{K}_{\exp}=\mathrm{cl}(\{\bm{\omega}\in\mathbb{R}^{3}:\omega_{1}\geq\omega_{2}\cdot\exp(\omega_{3}/\omega_{2}),\ \omega_{1}>0,\ \omega_{2}>0\})$ where $\mathrm{cl}$ is the closure operator. The Lipschitz modulus of a univariate function $f$ is defined as $\mathrm{Lip}(f):=\sup_{z,z^{\prime}\in\mathbb{R}}\left\{\lvert f(z)-f(z^{\prime})\rvert/\lvert z-z^{\prime}\rvert:z\neq z\right\}$ whereas its effective domain is $\mathrm{dom}(f)=\{z:f(z)<+\infty\}$. For a function $f:\mathbb{R}^{n}\mapsto\mathbb{R}$, its convex conjugate is $f^{*}(\bm{z})=\sup_{\bm{x}\in\mathbb{R}^{n}}\bm{z}^{\top}\bm{x}-f(\bm{x})$. We reserve $\alpha\geq 0$ for the radii of the norms of adversarial attacks on the features and $\varepsilon\geq 0$ for the radii of distributional ambiguity sets.
Appendix B MISSING PROOFS
B.1 Proof of Observation 1
@@
For any $\bm{\beta}\in\mathbb{R}^{n}$, with standard robust optimization arguments (Ben-Tal et al., 2009; Bertsimas and Den Hertog, 2022), we can show that
@@
where the first step follows from the definition of logloss, the second step follows from the fact that $\log$ and $\exp$ are increasing functions, the third step takes the constant terms out of the $\sup$ problem and exploits the fact that the optimal solution of maximizing a linear function will be at an extreme point of the $\ell_{p}$-ball, the fourth step uses the definition of dual norm, and finally, the redundant $-y\in\{-1,+1\}$ is omitted from the dual norm. We can therefore define the adversarial loss $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y):=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ where $\alpha$ models the strength of the adversary, $\bm{\beta}$ is the decision vector, and $(\bm{x},y)$ is an instance. Replacing ${\sup_{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}$ in DR-ARO with $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$ concludes the equivalence of the optimization problem.
@@ B.2 Proof of Corollary 1
Since the univariate loss $L^{\alpha}(z):=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ satisfying the identity $L^{\alpha}(\langle y\cdot\bm{x},\bm{\beta}\rangle)=\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$ is Lipschitz continuous, Theorem 14 (ii) of Shafieezadeh-Abadeh et al. (2019) is immediately applicable. We can therefore rewrite (29) as:
@@
B.3 Proof of Proposition 1
We prove Proposition 1 by constructing the optimization problem in its statement. We will thus dualize the inner $\sup$-problem of Inter-ARO for fixed $\bm{\beta}$. To this end, we present a sequence of reformulations to the inner problem and then exploit strong semi-infinite duality.
@@ B.3 Proof of Proposition 1
The following semi-infinite optimization problem, obtained by standard algebraic duality, is a strong dual to the above problem since $\varepsilon,\widehat{\varepsilon}>0$ (Shapiro, 2001).
@@
$\begin{array}[]{cll}\displaystyle\underset{\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}}}{\mathrm{minimize}}&\displaystyle\dfrac{1}{N}\left[N\varepsilon\lambda+\widehat{N}\widehat{\varepsilon}\widehat{\lambda}+\sum_{i=1}^{N}s_{i}+\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\right]&\\[14.22636pt] \mathrm{subject\;to}&\displaystyle\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\{\ell^{\alpha}_{\bm{\beta}}(\bm{x},+1)-\lambda\lVert\bm{x}^{i}-\bm{x}\rVert_{q}-\dfrac{\widehat{N}}{N}\widehat{\lambda}\lVert\widehat{\bm{x}}^{j}-\bm{x}\rVert_{q}\}\leq\\[8.53581pt] &{\quad s_{i}+\kappa\dfrac{1-y^{i}}{2}\lambda+\dfrac{\widehat{N}}{N}\cdot\left[\widehat{s}_{j}+\kappa\dfrac{1-\widehat{y}^{j}}{2}\widehat{\lambda}\right]}&\forall i\in[N],\ \forall j\in[\widehat{N}]\\[14.22636pt] &\displaystyle\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\{\ell^{\alpha}_{\bm{\beta}}(\bm{x},-1)-\lambda\lVert\bm{x}^{i}-\bm{x}\rVert_{q}-\dfrac{\widehat{N}}{N}\widehat{\lambda}\lVert\widehat{\bm{x}}^{j}-\bm{x}\rVert_{q}\}\leq\\[8.53581pt] &{\quad s_{i}+\kappa\dfrac{1+y^{i}}{2}\lambda+\dfrac{\widehat{N}}{N}\cdot\left[\widehat{s}_{j}+\kappa\dfrac{1+\widehat{y}^{j}}{2}\widehat{\lambda}\right]}&\forall i\in[N],\ \forall j\in[\widehat{N}]\\[14.22636pt] &\lambda\geq 0,\ \widehat{\lambda}\geq 0,\ \bm{s}\in\mathbb{R}^{N}_{+},\ \widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+},\end{array}$
@@
$\begin{array}[]{lll}&\displaystyle\int_{\bm{x}\in\mathbb{R}^{n}}\mathbb{Q}_{+1}^{i}(\mathop{}\!\mathrm{d}\bm{x})+\mathbb{Q}_{-1}^{i}(\mathop{}\!\mathrm{d}\bm{x})\leq 1&\forall i\in[N],\\[14.22636pt] &\displaystyle\int_{\bm{x}\in\mathbb{R}^{n}}\widehat{\mathbb{Q}}_{+1}^{j}(\mathop{}\!\mathrm{d}\bm{x})+\widehat{\mathbb{Q}}_{-1}^{j}(\mathop{}\!\mathrm{d}\bm{x})\leq 1&\forall j\in[\widehat{N}],\end{array}$
@@
Combining all the $\sup$-constraints with the help of an an auxiliary parameter $l\in\{-1,1\}$ and replacing this problem with the inner problem of Inter-ARO concludes the proof.∎
B.4 Proof of Proposition 2
We first present a technical lemma that will allow us to rewrite a specific type of difference of convex functions (DC) maximization problem that appears in the constraints of Inter-ARO. Rewriting such DC maximization problems is one of the key steps in reformulating Wasserstein DRO problems, and our lemma is inspired from Shafieezadeh-Abadeh et al. (2019, Lemma 47)Shafieezadeh-Abadeh et al. (2023, Theorem 3.8), and Belbasi et al. (2023, Lemma 1) who reformulate maximizing the difference of a convex function and a norm. Our DRO problem Inter-ARO, however, comprises two ambiguity sets, hence the DC term that we investigate will be the difference between a convex function and the sum of two norms. This requires a new analysis and we will see that Inter-ARO is NP-hard due to this additional difficulty.